OpenStudy (anonymous):

Find the maximum y-value on the graph of y=f(x) f(x)= -x^2+4x+9

OpenStudy (bahrom7893):

okay finally some calc1

OpenStudy (anonymous):

i've got a couple of questions too!

OpenStudy (bahrom7893):

y = -x^2 + 4x + 9 to maximize that take the derivative and set that equal to 0

OpenStudy (bahrom7893):

y' = -2x + 4 = 0 -2x + 4 = 0 -2x = -4 x = 2

OpenStudy (bahrom7893):

Soo when x = 2, y = -2^2 + 4*2 + 9 = -4 + 8 + 9 = 4+9 = 13

OpenStudy (anonymous):

Would or could you help with more of my questions?

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

sure go ahead, ask away bud

OpenStudy (bahrom7893):

make me feel like i actually know something lol

OpenStudy (anonymous):

for y=x^2-1 do the following: a. Sketch a graph of the equation b> identify the vertex compare the graph of y=f(x) to the graph of y=x^2 state any transformation used

OpenStudy (bahrom7893):

it's just a parabola with a vertex of (0,-1), pointing upward. Transformation is a shift down along the y axis i guess, not sure what the terms are..

OpenStudy (anonymous):

so when graphing what points so I use with the parabola.

OpenStudy (anonymous):

transformations would be down, left, right or up?

OpenStudy (anonymous):

Rather than using Calculus, we can use basic Pre-Calculus mathematics in order to find the maximum point. We know that the leading coefficient is negative, hence the vertex of the form must be a maximum. We can find this coordinate for the maximum by Completing-the-square and re-writing the function in vertex form. \[-x^2+4x+9=-(x-2)^2+13\] This means that the coordinate of the vertex (maximum in this case) is (2, 13). @PECKCE

OpenStudy (bahrom7893):

transformation is down

OpenStudy (bahrom7893):

ok let's generate a bunch of points for you to use: (0,-1) (1,0);(-1,0) (2,3);(-2;3), etc.

OpenStudy (anonymous):

Which means the maximum y-value is y = 13, and it occurs at x = 2. Do you understand? I'm hoping you're familiar with completing the square/vertex form. @PECKCE

OpenStudy (anonymous):

thank you @genius12 it was very kind of you, I am not familiar with that but @bahrom7893 explain it to me and came to the same solution.

OpenStudy (anonymous):

@PECKCE What @bahrom7893 has done is completely valid but it requires you to know Calculus. If you haven't covered Calculus, the approach would seem difficult to understand. And you should be familiar with completing the square/vertex form, if you're not, go here to get a tutorial: http://www.mathsisfun.com/algebra/completing-square.html That will teach you exactly what to do and what completing the square means.